b. In a quadrilateral, the sum of the measures of all the four angles is 360o. Suppose a quadrilateral has two angles that are equal. Also, the third angle is equal to the sum of the two equal angles. The fourth angle is 60o less than twice the sum of the other three angles. Find the measures of the angles in the quadrilateral.
Answer by macston(5194)

.
f=first angle; s=second angle=f; t=third angle=f+s=2f;
g=fourth angle=2(f+s+t)-60=2(f+f+2f)-60=2(4f)-60=8f-60
.
f+s+t+g=360 degrees Substitute for s,t and g.
f+f+2f+8f-60=360 degrees
12f=420 degrees
f=35 degrees ANSWER 1: First angle is 35 degrees.
s=f=35 degrees ANSWER 2: Second angle is 35 degrees.
t=f+s=35+35=70 degrees ANSWER 3: Third angle is 70 degrees.
g=8f=8(35)-60=220 degrees ANSWER 4: Fourth angle is 220 degrees.